# On the inclusion of non-concurrent controls in platform trials with a futility interim analysis

### v1

@@ 4.1 Mean adjusted estimator
To account for the bias due to the interim analysis for Arm 1, we consider a mean adjusted estimator (MAE) [22], constructed by subtracting the estimated conditional bias of $\tilde{\theta}_{2}$ (conditional on the event that Arm 1 continues after the interim analysis) $E[\tilde{\theta}_{2}-\theta_{2}|Z_{11}\geq c_{1}]$ from the original model-based estimator, $\tilde{\theta}_{2}$, in cases when Arm 1 continues after the interim analysis. The resulting bias-adjusted estimator of the treatment effect in Arm 2, $\tilde{\theta}^{A}_{2}$, is given by:
$\tilde{\theta}^{A}_{2}=\begin{cases}\bar{y}_{22}-\bar{y}_{02},&\text{if }Z_{11}<c_{1},\\[6.0pt] \tilde{\theta}_{2}-\widehat{B}\left[\tilde{\theta}_{2}\mid Z_{11}\geq c_{1}\right],&\text{if }Z_{11}\geq c_{1},\end{cases}$ (8)
@@
$\widehat{B}\left[\tilde{\theta}_{2}\mid Z_{11}\geq c_{1}\right]=\frac{\varrho\cdot\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}\cdot\phi(\hat{\gamma})}{1-\Phi(\hat{\gamma})}\text{, \hskip 5.69046pt where \hskip 5.69046pt }\hat{\gamma}=c_{1}-\frac{\hat{\theta}_{1}}{\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}}$ (9)
The term $\widehat{B}[\tilde{\theta}_{2}|Z_{11}\geq c_{1}]$ denotes the estimator of the conditional bias defined in (6). Note that we use the conditional bias of $\tilde{\theta}_{2}$ for the adjustment, because the MAE is only employed in case Arm 1 continues, as there is no bias in the effect estimation for Arm 2 if Arm 1 stops at the interim.
The expression for the conditional bias (6) depends on the futility bound and the standardized treatment effect of Arm 1. While the sample sizes, futility bound, and sample variance are considered as known parameters here, the treatment effect in Arm 1, $\theta_{1}$, is an unknown parameter. However, given that the mean adjusted estimator is computed at the end of the trial, we can estimate $\theta_{1}$ using the observed data. Therefore, in order to estimate $\widehat{B}[\tilde{\theta}_{2}|Z_{11}\geq c_{1}]$, we can use a plug-in estimator for $\theta_{1}$, denoted by $\hat{\theta}_{1}$. In particular, we consider four approaches for estimating $\theta_{1}$: (i) using the sample means of responses from Arm 1 and the control arm from the whole trial, denoted by $\bar{y}_{1\cdot}$ and $\bar{y}_{0\cdot}$ respectively, that is $\hat{\theta}_{1}=\bar{y}_{1\cdot}-\bar{y}_{0\cdot}$; (ii) using only the sample means of responses from period 1, that is $\hat{\theta}_{1}=\bar{y}_{11}-\bar{y}_{01}$; (iii) using only data from period 2, that is $\hat{\theta}_{1}=\bar{y}_{12}-\bar{y}_{02}$; and (iv) the conditional uniform minimum variance unbiased estimator (CUMVUE), denoted by $\hat{\theta}_{1}^{CUMVUE}$, derived by applying the Rao-Blackwell theorem. The estimators (iii) and (iv) are unbiased estimators of $\theta_{1}$ conditionally on continuing after the interim.
@@
$\hat{\theta}_{1}^{UMVUE}=(\bar{y}_{1\cdot}-\bar{y}_{0\cdot})-\frac{I_{2}-I_{1}}{I_{2}\sqrt{I_{1}}}\cdot\frac{-\phi(c_{1},Z_{12}\sqrt{I_{1}/I_{2}},(I_{2}-I_{1})/I_{2})}{1-\Phi(c_{1},Z_{12}\sqrt{I_{1}/I_{2}},(I_{2}-I_{1})/I_{2})}$ (10)
where $Z_{12}=(\bar{y}_{1\cdot}-\bar{y}_{0\cdot})/\sigma\sqrt{1/(n_{11}+n_{12})+1/(n_{01}+n_{02})}$ denotes the Z-statistic for arm 1 at the end of the trial, and $I_{1}=1/(\sigma^{2}(1/n_{11}+1/n_{01}))$ and $I_{2}=1/(\sigma^{2}(1/(n_{11}+n_{12})+1/(n_{01}+n_{02})))$ are the information levels in periods 1 and 2, respectively. The CUMVUE for $\theta_{1}$ after period 2 is then constructed using (10) as follows:
$\hat{\theta}_{1}^{CUMVUE}=\frac{Z_{12}\cdot\sqrt{I_{2}}-I_{1}\cdot\hat{\theta}_{1}^{UMVUE}}{(I_{2}-I_{1})}$ (11)
@@ 4.1.1 Hypothesis testing
$T=\begin{cases}\tilde{\theta}^{A}_{2}/\sigma\sqrt{1/n_{02}+1/n_{22}},&\text{if }Z_{11}<c_{1},\\[6.0pt] \tilde{\theta}^{A}_{2}/\sqrt{\widehat{\text{Var}}(\tilde{\theta}^{A}_{2}\mid Z_{11}\geq c_{1})},&\text{if }Z_{11}\geq c_{1},\end{cases}$ (12)
Under the null hypothesis $H_{0,2}$, the test statistic $T$ follows the standard normal distribution. We reject $H_{0,2}$ at the one-sided significance level $\alpha$ if $T>\Phi^{-1}(1-\alpha)$. Note that for the case where Arm 1 stops at the interim analysis, this is equivalent to the standard z-test.
To obtain an estimator of the variance of the mean adjusted estimator $\tilde{\theta}^{A}_{2}$ in cases where Arm 1 continues, we employ the bootstrap. Specifically, the conditional variance $\text{Var}(\tilde{\theta}^{A}_{2}\mid Z_{11}\geq c_{1})$ is estimated via bootstrap stratified per arm and period. Letting $Y^{i*}_{ks}$ denote a response observation randomly drawn from arm $k$ and period $s$, we considered the following bootstrap algorithm:
@@
    Draw two independent bootstrap samples $Y^{1*}_{11},\ldots,Y^{n_{11}*}_{11}$ and $Y^{1*}_{01},\ldots,Y^{n_{01}*}_{01}$ from period 1 observations in Arm 1 and control, respectively.
@@
    If $Z_{11}^{*}<c_{1}$, discard this resample and restart at Step 1. Otherwise, continue to Step 4.
@@
    Draw two independent bootstrap samples $Y^{1*}_{12},\ldots,Y^{n_{12}*}_{12}$ and $Y^{1*}_{02},\ldots,Y^{n_{02}*}_{02}$ from period 2 of the trial; from Arm 1 and control, respectively.
@@
    $\widehat{\text{Var}}\left(\tilde{\theta}^{A}_{2}\mid Z_{11}\geq c_{1}\right)=\frac{1}{B}\sum_{j=1}^{B}\left(\tilde{\theta}_{2,j}^{A*}-\frac{1}{B}\sum_{j=1}^{B}\tilde{\theta}_{2,j}^{A*}\right)^{2}$
Note that the bootstrap procedure mimics the course of the original trial, preserving the sample sizes per arm and period, as well as the futility stopping rule used at the interim analysis. In order to base the variance estimation on an equal number of bootstrap estimates (denoted by $B$), regardless of the futility bound $\alpha_{1}$, we repeat the bootstrap procedure until we obtain $B$ samples where Arm 1 has continued after the interim analysis.
@@ 5.1 Evaluated operating characteristics
In the simulation study presented below, we therefore investigate the newly proposed estimator in terms of the bias, root mean squared error, and the corresponding hypothesis test in terms of type I error rate, and statistical power. We discuss the marginal and conditional bias of both considered methods. For the remaining operating characteristics, we focus on its value conditional on Arm 1 continuing after the interim analysis, since the mean adjusted estimator is only employed in this case. Results for marginal operating characteristics can be found in the supplementary material. We compare the performance of the proposed mean adjusted estimate to the unadjusted model-based estimate and the separate analysis, which only considers data from period 2. All simulation scenarios were replicated 100.000 times to estimate the investigated operating characteristics. In the bootstrap algorithm for estimating the variance of $\tilde{\theta}^{A}_{2}$, we used $B=1000$ replicates.
@@ 5.2 Data generation and considered design parameters
Consider a platform trial with two experimental arms, as described in Section 2 and illustrated in Figure 1. The continuous outcome $y_{j}$ for patient $j$ is then drawn from a normal distribution according to $y_{j}\sim\mathcal{N}(\mu,\sigma^{2})$ with
@@
Table 1 summarizes the values of the design parameters that we consider in the simulation study. In particular, the futility bound $\alpha_{1}$ varies from 0.1 to 0.95. We do not consider the extreme bounds of 0 and 1, as they result in either always conducting a separate analysis based only on the period 2 data or always using the regression model, and thus the analysis is marginally unbiased in these cases. The sample sizes per arm and period are varied in such a way that the impact of two ratios on the operating characteristics can be investigated. First, the ratio between the sample sizes in periods 1 and 2, $r=n_{01}/n_{02}=n_{11}/n_{12}$, which is varied from 1/15 to 10. Thus, the sample size per arm in period 1 can then be expressed as a multiple of the sample size per arm in period 2, which is kept constant. Second, the allocation ratio in Arm 1 and the control arm, denoted by $a=n_{11}/n_{01}=n_{12}/n_{02}$, is also varied from 1/15 to 10, allowing to examine the behavior of the methods under different allocation ratios. In particular, the sample size in Arm 1 is $a$-times larger/smaller with respect to the sample size in the control arm, which is kept constant. The sample sizes in Arm 2 and the control arm in period 2 are always set to 150. The strength of the time trend $\lambda$ is varied from 0.15-0.15 to $0.15$, i.e., considering both increase and decrease in the mean response over time.
@@
Futility bound $\alpha_{1}$ 0.1, 0.15, 0.2, 0.25, 0.35, 0.50, 0.65, 0.75, 0.95
@@
Table 1: Considered values of the design parameters varied in the simulations and the resulting sample sizes. When varying one design parameter, the remaining ones were kept constant at the value highlighted in bold type. In total, 28 scenarios were considered under both the null and alternative hypotheses.
We focus on testing the null hypothesis $H_{0,2}:\theta_{2}=\mu_{2}-\mu_{0}=0$ against the one-sided alternative $H_{1,2}:\theta_{2}>0$ at the significance level $\alpha=0.025$ using the proposed mean adjusted estimator $\tilde{\theta}^{A}_{2}$, as well as the unadjusted model-based estimator $\tilde{\theta}_{2}$. For comparison, we include results for the root mean squared error and statistical power obtained when using the separate analysis, regardless of the interim outcome. We consider scenarios where Arm 2 is under the null hypothesis ($\theta_{2}=0$), as well as cases where the alternative holds. In the latter case, the treatment effect for Arm 2 is set to $\theta_{2}=0.32$ such that the separate analysis comparing Arm 2 to the concurrent controls leads to approximately 80% power using the given sample sizes. In the simulations, we only consider cases where the null hypothesis holds for Arm 1 ($\theta_{1}=0$). This is sufficient, as the results for Arm 2 depend on the probability of Arm 1 continuing to period 2, which is based on the combination of both the treatment effect and the rule used for the interim analysis.

### v2

@@ 4.1 Mean adjusted estimator
To account for the bias due to the interim analysis for Arm 1, we consider a mean adjusted estimator (MAE) [32], constructed by subtracting the estimated conditional bias of $\tilde{\theta}_{2}$ (conditional on the event that Arm 1 continues after the interim analysis) $E[\tilde{\theta}_{2}-\theta_{2}|c_{F}\leq Z_{11}\leq c_{E}]$ from the original model-based estimator, $\tilde{\theta}_{2}$, in cases when Arm 1 continues after the interim analysis. The resulting bias-adjusted estimator of the treatment effect in Arm 2, $\tilde{\theta}^{A}_{2}$, is given by:
$\tilde{\theta}^{A}_{2}=\begin{cases}\bar{y}_{22}-\bar{y}_{02},&\text{if }Z_{11}<c_{F}\text{ or }Z_{11}>c_{E},\\[6.0pt] \tilde{\theta}_{2}-\widehat{B}\left[\tilde{\theta}_{2}\mid c_{F}\leq Z_{11}\leq c_{E}\right],&\text{if }c_{F}\leq Z_{11}\leq c_{E},\end{cases}$ (6)
@@
$\widehat{B}\left[\tilde{\theta}_{2}\mid c_{F}\leq Z_{11}\leq c_{E}\right]=\frac{\varrho\cdot\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}\cdot\left(\phi\left(c_{F}-\hat{\delta}\right)-\phi\left(c_{E}-\hat{\delta}\right)\right)}{\Phi\left(c_{E}-\hat{\delta}\right)-\Phi\left(c_{F}-\hat{\delta}\right)}\text{, \hskip 5.69046pt where \hskip 5.69046pt }\hat{\delta}=\frac{\hat{\theta}_{1}}{\sigma\sqrt{\frac{1}{n_{11}}+\frac{1}{n_{01}}}}$
The term $\widehat{B}[\tilde{\theta}_{2}|c_{F}\leq Z_{11}\leq c_{E}]$ denotes the estimator of the conditional bias defined in (4). Note that we use the conditional bias of $\tilde{\theta}_{2}$ for the adjustment, because the MAE is only employed in case Arm 1 continues, as there is no bias in the effect estimation for Arm 2 if Arm 1 stops at the interim.
The expression for the conditional bias (4) depends on the futility bound and the standardized treatment effect of Arm 1. While the sample sizes, futility bound, and sample variance are considered as known parameters here, the treatment effect in Arm 1, $\theta_{1}$, is an unknown parameter. However, given that the mean adjusted estimator is computed at the end of the trial, we can estimate $\theta_{1}$ using the observed data. Therefore, in order to estimate $\widehat{B}[\tilde{\theta}_{2}|c_{F}\leq Z_{11}\leq c_{E}]$, we can use a plug-in estimator for $\theta_{1}$, denoted by $\hat{\theta}_{1}$. In particular, we consider four approaches for estimating $\theta_{1}$: (i) using the sample means of responses from Arm 1 and the control arm from the whole trial, denoted by $\bar{y}_{1\cdot}$ and $\bar{y}_{0\cdot}$ respectively, that is $\hat{\theta}_{1}=\bar{y}_{1\cdot}-\bar{y}_{0\cdot}$; (ii) using only the sample means of responses from period 1, that is $\hat{\theta}_{1}=\bar{y}_{11}-\bar{y}_{01}$; (iii) using only data from period 2, that is $\hat{\theta}_{1}=\bar{y}_{12}-\bar{y}_{02}$; and (iv) the conditional uniform minimum variance unbiased estimator (CUMVUE), denoted by $\hat{\theta}_{1}^{CUMVUE}$, derived by applying the Rao-Blackwell theorem. The estimators (iii) and (iv) are unbiased estimators of $\theta_{1}$ conditionally on continuing after the interim.
@@
$\hat{\theta}_{1}^{UMVUE}=(\bar{y}_{1\cdot}-\bar{y}_{0\cdot})-\frac{I_{2}-I_{1}}{I_{2}\sqrt{I_{1}}}\cdot\frac{\phi(c_{E},Z_{12}\sqrt{I_{1}/I_{2}},(I_{2}-I_{1})/I_{2})-\phi(c_{F},Z_{12}\sqrt{I_{1}/I_{2}},(I_{2}-I_{1})/I_{2})}{\Phi(c_{E},Z_{12}\sqrt{I_{1}/I_{2}},(I_{2}-I_{1})/I_{2})-\Phi(c_{F},Z_{12}\sqrt{I_{1}/I_{2}},(I_{2}-I_{1})/I_{2})}$ (7)
where $Z_{12}=(\bar{y}_{1\cdot}-\bar{y}_{0\cdot})/\sigma\sqrt{1/(n_{11}+n_{12})+1/(n_{01}+n_{02})}$ denotes the Z-statistic for arm 1 at the end of the trial, and $I_{1}=1/(\sigma^{2}(1/n_{11}+1/n_{01}))$ and $I_{2}=1/(\sigma^{2}(1/(n_{11}+n_{12})+1/(n_{01}+n_{02})))$ are the information levels in periods 1 and 2, respectively. $\phi(x,\mu,\sigma^{2})$ and $\Phi(x,\mu,\sigma^{2})$ are the probability density and cumulative distribution functions of a normal distribution with mean $\mu$ and variance $\sigma^{2}$ evaluated at $x$, respectively. The CUMVUE for $\theta_{1}$ after period 2 is then constructed using (7) as follows:
$\hat{\theta}_{1}^{CUMVUE}=\frac{Z_{12}\cdot\sqrt{I_{2}}-I_{1}\cdot\hat{\theta}_{1}^{UMVUE}}{(I_{2}-I_{1})}$
@@ 4.1.1 Hypothesis testing
$T=\begin{cases}\tilde{\theta}^{A}_{2}/\sigma\sqrt{1/n_{02}+1/n_{22}},&\text{if }Z_{11}<c_{F}\text{ or }Z_{11}>c_{E},\\[6.0pt] \tilde{\theta}^{A}_{2}/\sqrt{\widehat{\text{Var}}(\tilde{\theta}^{A}_{2}\mid c_{F}\leq Z_{11}\leq c_{E}}),&\text{if }c_{F}\leq Z_{11}\leq c_{E},\end{cases}$
Under the null hypothesis $H_{0,2}$, the test statistic $T$ approximately follows the standard normal distribution. We reject $H_{0,2}$ at the one-sided significance level $\alpha$ if $T>\Phi^{-1}(1-\alpha)$. Note that for the case where Arm 1 stops at the interim analysis, this is equivalent to the standard z-test.
To obtain an estimator of the variance of the mean adjusted estimator $\tilde{\theta}^{A}_{2}$ in cases where Arm 1 continues, we employ the bootstrap. Specifically, the conditional variance $\text{Var}(\tilde{\theta}^{A}_{2}\mid c_{F}\leq Z_{11}\leq c_{E})$ is estimated via bootstrap stratified per arm and period. Letting $Y^{i*}_{ks}$ denote a response observation randomly drawn from arm $k$ and period $s$, we considered the following bootstrap algorithm:
@@
    Draw two independent bootstrap samples $Y^{1*}_{01},\ldots,Y^{n_{01}*}_{01}$ and $Y^{1*}_{11},\ldots,Y^{n_{11}*}_{11}$ from period 1 observations in control and Arm 1, respectively.
@@
    If $Z_{11}^{*}<c_{F}$ or $Z_{11}^{*}>c_{E}$, discard this resample and restart at Step 1. Otherwise, continue to Step 4.
@@
    Draw three independent bootstrap samples $Y^{1*}_{02},\ldots,Y^{n_{02}*}_{02}$, $Y^{1*}_{12},\ldots,Y^{n_{12}*}_{12}$, and $Y^{1*}_{22},\ldots,Y^{n_{22}*}_{22}$ from period 2 of the trial; from control, Arm 1, and Arm 2, respectively.
@@
    $\widehat{\text{Var}}\left(\tilde{\theta}^{A}_{2}\mid c_{F}\leq Z_{11}\leq c_{E}\right)=\frac{1}{B}\sum_{j=1}^{B}\left(\tilde{\theta}_{2,j}^{A*}-\frac{1}{B}\sum_{j=1}^{B}\tilde{\theta}_{2,j}^{A*}\right)^{2}$
Note that the bootstrap procedure mimics the course of the original trial, preserving the sample sizes per arm and period, as well as the futility and efficacy stopping rules used at the interim analysis. In order to base the variance estimation on an equal number of bootstrap estimates (denoted by $B$), regardless of the futility and efficacy bounds $\alpha_{F}$ and $\alpha_{E}$, we repeat the bootstrap procedure until we obtain $B$ samples where Arm 1 has continued after the interim analysis.
@@ 5.1 Evaluated operating characteristics
In the simulation study presented below, we therefore investigate the newly proposed estimator in terms of the bias and root mean squared error, and the corresponding hypothesis test in terms of type I error rate and statistical power. We discuss the marginal and conditional bias of both considered methods. For the remaining operating characteristics, we focus on their value conditional on Arm 1 continuing after the interim analysis, since the mean adjusted estimator is only employed in this case. Results for marginal operating characteristics can be found in the supplementary material (Section A). We compare the performance of the proposed mean adjusted estimate to the unadjusted model-based estimate and the separate analysis, which only considers data from period 2. All simulation scenarios were replicated 100.000 times to estimate the investigated operating characteristics. In the bootstrap algorithm for estimating the variance of $\tilde{\theta}^{A}_{2}$, we used $B=1000$ replicates.
@@ 5.2 Data generation and considered design parameters
Consider a platform trial with two experimental arms, as described in Section 2 and illustrated in Figure 1. We assume uniform recruitment of one patient per time unit, such that $t_{j}=j$, and block randomization with block sizes of 4 and 6 in periods 1 and 2, respectively. The continuous outcome $y_{j}$ for patient $j$ is then drawn from a normal distribution according to $y_{j}\sim\mathcal{N}(\mu,\sigma^{2})$ with
@@
  • Seasonal time trend: $f(j)=\lambda\cdot\mathrm{sin}\big(2\pi\cdot\frac{j-1}{N-1}\big)$, where $N$ indicates the total sample size in the trial. The seasonal trend consists of one cycle, where the mean response increases at first and decreases afterwards, while the respective peaks of the cycle correspond to $\lambda$ or $\lambda$.
  • Time trend following a random walk: $\{f(j)\}_{j=1}^{N}$, defined by $f(1)=0$ and $f(j)=f(j-1)+\frac{\lambda}{N-1}\cdot\varepsilon_{j}$ for $j=2,\ldots,N$, where $\text{I\kern-1.49994ptP}(\varepsilon_{j}=1)=\text{I\kern-1.49994ptP}(\varepsilon_{j}=-1)=0.5$. Hence, the mean response for patient $j$ increases or decreases by $\frac{\lambda}{N-1}$ from the mean response of patient j1j-1, with probability 0.5, respectively. For patient $j=1$, no time trend is assumed.
Table 1 summarizes the values of the design parameters that we consider in the simulation study. In particular, the treatment effect in Arm 1 $\theta_{1}$ varies from -1 to 1. The sample sizes per arm and period are varied in such a way that the impact of two ratios on the operating characteristics can be investigated. First, the ratio between the sample sizes in periods 1 and 2, $r=n_{01}/n_{02}=n_{11}/n_{12}$, which is varied from 1/15 to 10. Thus, the sample size per arm in period 1 can then be expressed as a multiple of the sample size per arm in period 2, which is kept constant. Second, the allocation ratio in Arm 1 and the control arm, denoted by $a=n_{11}/n_{01}=n_{12}/n_{02}$, is also varied from 1/15 to 10, allowing to examine the behavior of the methods under different allocation ratios. In particular, the sample size in Arm 1 is larger/smaller by a factor of $a$ with respect to the sample size in the control arm, which is kept constant. The sample sizes in Arm 2 and the control arm in period 2 are always set to 150. The strength of the time trend $\lambda$ is varied from 0.15-0.15 to $0.15$, i.e., considering both increase and decrease in the mean response over time. In settings with varying $r$, $a$, and $\lambda$, the treatment effect $\theta_{1}$ was chosen such that a fixed design analysis of Arm 1 would result in approximately 80% power.
@@
Treatment effect $\theta_{1}$
[-1, 1]; with increments of 0.2 in [-1, -0.2] and [0.6, 1]
and increments of 0.05 in [-0.15, 0.5],
resulting in 22 values in total
@@
Table 1: Considered values of the design parameters varied in the simulations and the resulting sample sizes. When varying one design parameter, the remaining ones were kept constant at the value highlighted in bold type. In each case, the futility boundary was set to $\alpha_{F}=0.5$, while the efficacy boundary $\alpha_{E}$ was set to the O’Brien-Fleming boundary for the given sample sizes and $\alpha_{F}$. In settings with varying $r$, $a$, and $\lambda$, the treatment effect $\theta_{1}$ was chosen such that a fixed design analysis of Arm 1 would have 80% power. In total, 41 scenarios were considered under both the null and alternative hypotheses. Note that not all combinations are considered.
We focus on testing the null hypothesis $H_{0,2}:\theta_{2}=\mu_{2}-\mu_{0}=0$ against the one-sided alternative $H_{1,2}:\theta_{2}>0$ at the significance level $\alpha=0.025$ using the proposed mean adjusted estimator $\tilde{\theta}^{A}_{2}$, as well as the unadjusted model-based estimator $\tilde{\theta}_{2}$. For comparison, we include results for the root mean squared error and statistical power obtained when using the separate analysis, regardless of the interim outcome. We consider scenarios where Arm 2 is under the null hypothesis ($\theta_{2}=0$), as well as cases where the alternative holds. In the latter case, the treatment effect for Arm 2 is set to $\theta_{2}=0.32$ such that the separate analysis comparing Arm 2 to the concurrent controls leads to approximately 80% power using the given sample sizes.